SUPPOSE EACH ORDER PAIRS ARE FOR THE SAME DIRECT VARIATION. FIND EACH MISSING VALUE

A. (-1,2) AND (4,Y)
B. (X,-4) AND (-9,-12)
PLEASE HELP ME!!!!!!!!!!!!!!

Answer by stanbon(75887)

A. (-1,2) AND (4,Y)
The ratio of 2/(-1)=-2
Then Y/4=-2
Y=-8
B. (X,-4) AND (-9,-12)
The ratio -12/-9= 4/3
Then -4/X=4/3
And -1/x=1/3
X=-3
Cheers,
Stan H.